# Towards Mitigating Modality Bias in Vision-Language Models for Temporal Action Localization

### v1

@@ Abstract
Temporal Action Localization (TAL) requires identifying both the boundaries and categories of actions in untrimmed videos. While vision-language models (VLMs) offer rich semantics to complement visual evidence, existing approaches tend to overemphasize linguistic priors at the expense of visual performance, leading to a pronounced modality bias. We propose ActionVLM, a vision-language aggregation framework that systematically mitigates modality bias in TAL. Our key insight is to preserve vision as the dominant signal while adaptively exploiting language only when beneficial. To this end, we introduce (i) a debiasing reweighting module that estimates the language advantage—the incremental benefit of language over vision-only predictions—and dynamically reweights language modality accordingly, and (ii) a residual aggregation strategy that treats language as a complementary refinement rather than the primary driver. This combination alleviates modality bias, reduces overconfidence from linguistic priors, and strengthens temporal reasoning. Experiments on THUMOS14 show that our model outperforms state-of-the-art by up to 3.2% mAP.
@@
Jiaqi Li1, Guangming Wang2, Shuntian Zheng1, Minzhe Ni1, Xiaoman Lu1, Guanghui Ye3, Yu Guan1 1 UVLab, Department of Computer Science, University of Warwick 2 Department of Automation, University of Cambridge 3 College of Computer Science and Electronic Engineering, Hunan University
@@ 2.3 Modality Bias
Prior work can be grouped into two directions. Diagnostic works primarily focus on revealing bias. For example, gradient-based analyses Kwon et al. (2025) and the attention sink studies Xiao et al. (2023) provide valuable insights into how linguistic shortcuts dominate inference but provide little in terms of mitigation. In contrast, mitigation methods typically rely on either extra branches Cadene et al. (2019); Clark et al. (2019), which introduce a language-only predictor to discount bias but nearly double computation, or counterfactual data augmentations Chen et al. (2020), which enforce visual reliance through handcrafted or synthetic mismatched pairs. Although effective, both incur high costs or poor scalability, limiting their applicability to large-scale video understanding.
@@ 3.3 Language Advantage
We introduce Language Advantage (LA), denoted as $A_{lang}$, to guide the VLM in dynamically adjusting the language modality weight for each frame. Conceptually, LA functions as a lightweight estimate of information gain: when language helps disambiguate visual patterns, its advantage increases; when language introduces noise or redundancy, the advantage becomes small or negative, prompting the model to rely more on visual evidence.
Specifically for TAL, we formulate LA as: $A_{lang}=\mathcal{L}_{v}-\mathcal{L}_{vl}$, where $\mathcal{L}_{v}$ is the task loss computed on the vision-only branch, and $\mathcal{L}_{vl}$ the loss when language is also incorporated. A positive $A_{lang}$ indicates that language improves the prediction, while negative values imply redundancy or harmful bias of language modality.
Alternating-Epoch Estimation. A naïve implementation would require an additional vision-only branch to compute $\mathcal{L}_{v}$. To avoid this overhead, we adopt a lightweight alternating-epoch estimation inspired by reinforcement learning Schulman et al. (2015). Training alternates between vision-only and vision–language epochs. Vision-only epochs compute class-wise average losses $\bar{\mathcal{L}}_{v}^{(c)}$, which then serve as a stable proxy for $\mathcal{L}_{v}$ during vision-language epochs. The language advantage for the $l^{\text{th}}$ frame is thus estimated as
@@
This estimation scheme offers three practical benefits. 1) Lightweight computation. Both epochs share the same backbone and differ only by activating or skipping the language model, resulting in negligible overhead and no architectural duplication. 2) Efficient multimodal training. Vision-only epochs skip the language model entirely, making their computation cheaper than standard multimodal updates, while still providing a principled signal for calibrating modality contributions during TAL optimization. 3) Model and dataset agnosticism. Because LA is derived solely from task losses rather than dataset-level statistics, the estimation remains stable across different datasets and compatible with a variety of VLM backbones.
@@ 6   Backpropagate vision-only task loss: $\mathcal{L}_{v}$;
Algorithm 1 outlines our LA estimation, where the vision-only loss $\mathcal{L}_{v}$ is approximated by the class-wise average loss $\bar{\mathcal{L}}_{v}^{(c)}$ computed in the preceding vision-only epoch.
@@ 3.4 Debiasing Feature Aggregation
To dynamically regulate the contribution of the language modality, we derive a reliability weight $\lambda_{lang}$ from the predicted LA $\hat{A}_{lang}$. To avoid assigning positive weights to negative advantages, we clip it and scale it into $[0,1)$ through sigmoid mapping:
@@
Rather than relying solely on language features for downstream tasks, we preserve the visual feature $F_{vis}$ and treat language features as residual refinement. Inspired by residual learning He et al. (2016), this design allows language to complement vision only when beneficial, thereby mitigating modality bias. As illustrated in Fig. 3(b), the final representations for classification and localization are obtained via element-wise addition:
@@ 4.2 Comparison with SoTA Methods
To examine robustness to language model capacity, we evaluate two InternVL3 Chen et al. (2024b) variants that use Qwen2.5-1.5B and 3B. InternVL3 is chosen for its strong alignment with visual downstream tasks. ActionVLM-S-3B achieves a 0.6% mAP improvement over the 1.5B counterpart on THUMOS14, suggesting that broader knowledge and richer pretraining better support TAL.
@@ 4.3 Analysis on Modality Bias
Measuring Modality Bias. To quantify the reliance of the model on linguistic cues, we adopt Language Advantage on Performance (LAP), inspired by the diagnostic setup in Cadene et al. (2019). Unlike the loss-based language advantage used for training, LAP measures modality bias at the performance level. Specifically, we deliberately inject conflicting language priors (i.e., unmatched action descriptions) and record the relative drop in average mAP compared with the aligned setup. If the model is overly dependent on text, these contradictory cues will mislead prediction and yield a sharp mAP degradation, resulting in a large LAP. Conversely, a small drop indicates that the model relies primarily on visual evidence and remains robust under linguistic perturbations.
In Table 2, we also examine two forms of extreme modality-bias hallucinations: 1) Fixed outputs with obviously fake boundaries regardless of input. 2) Infinite listing or repeating segments.
@@ 4.4 Analysis on Language Advantage
To analyze how Language Advantage (LA) affects performance, we group action categories in THUMOS14 into easy, medium, and hard subsets based on the performance of the vision-only baseline (detailed in Appx. A.8). The hard subset includes Billiard and FrisbeeCatch. While multiple factors may contribute to their lower performance, both actions exhibit notable visual ambiguity: Billiard exhibits subtle and sparse motion cues, while FrisbeeCatch is easily confused with throwing actions and differs only in temporal pose order (Fig. 4).
@@ 5 Conclusion
In this study, we present ActionVLM, a Vision-Language Model (VLM) tailored for Temporal Action Localization (TAL). While conventional VLMs suffer from modality bias in vision-centric tasks, our feature aggregation formulation preserves visual grounding and selectively incorporates complementary language cues through language advantage estimation and debiasing residual reweighting. Extensive experiments show that ActionVLM effectively alleviates modality bias and enables language to serve as a reliable source of semantic refinement. Consequently, ActionVLM achieves consistent gains across TAL benchmarks, with notable improvements under visual ambiguity where prior methods struggle. Given its simplicity and negligible overhead, our debiasing formulation offers a practical path toward robust, vision-grounded multimodal learning.
Limitation
@@
This work builds upon publicly available video datasets and does not involve private or personally identifiable data. We acknowledge that vision-language models may inherit biases from pre-training corpora, and our method is intended solely for academic research to advance robust and fair video understanding.
@@ A.2 Analysis on Efficiency
Table 6 compares ActionVLM with both conventional VLMs and strong vision-only baselines in terms of computational efficiency and accuracy on THUMOS14. We report training and inference latency per video clip under the same configuration. The window size denotes the number of frames processed in a single temporal chunk, which directly affects both efficiency and temporal context coverage. We set a smaller window size for conventional VLMs to fit GPU memory, while ActionVLM and AdaTAD use longer windows to capture richer temporal dynamics following the original setting Liu et al. (2024).
@@ A.3 Additional Experiments on Epic-Kitchens
In order to validate the generalizability of our model on diverse datasets, we further conduct evaluation on EPIC-Kitchens 100 Damen et al. (2020). As an egocentric dataset collected from 700 egocentric daily kitchen activities, EPIC-Kitchens 100 exhibits fine-grained, long-tail action distributions and strong domain specificity that diverges from the visual and linguistic patterns commonly found in large-scale VLM pretraining corpora. Consequently, achieving high absolute mAP on EPIC-Kitchens is considerably more difficult Liu et al. (2024).
@@ A.6 Histogram of Linguistically Preferred Boundaries
A.7 Analysis on Performance Gain
@@
A.8 Action Categories by Visual Ambiguity

### v2

@@ Abstract
Temporal Action Localization (TAL) requires identifying both the boundaries and categories of actions in untrimmed videos. While vision-language models (VLMs) offer rich semantics to complement visual evidence, existing approaches tend to overemphasize linguistic priors at the expense of visual performance, leading to a pronounced modality bias. We propose ActionVLM, a vision-language aggregation framework that systematically mitigates modality bias in TAL. Our key insight is to preserve vision as the dominant signal while adaptively exploiting language only when beneficial. To this end, we introduce (i) a debiasing reweighting module that estimates the language advantage—the incremental benefit of language over vision-only predictions—and dynamically reweights language modality accordingly, and (ii) a residual aggregation strategy that treats language as a complementary refinement rather than the primary driver. This combination alleviates modality bias, reduces overconfidence from linguistic priors, and strengthens temporal reasoning. Experiments on THUMOS14 show that our model outperforms state-of-the-art by up to 3.2% mAP. Our code is available at https://github.com/JiaqiLi404/ActionVLM
@@
Jiaqi Li1, Guangming Wang2, Shuntian Zheng1, Minzhe Ni1, Xiaoman Lu1, Guanghui Ye3, Yu Guan1🖂thanks: Yu Guan is the corresponding author. 1 UVLab, Department of Computer Science, University of Warwick 2 Department of Automation, University of Cambridge 3 Department of Computer Science, The University of Sheffield {Jiaqi.Li.16, Shuntian.Zheng, Minzhe.Ni, Xiaoman.Lu, Yu.Guan}@warwick.ac.uk, gw462@cam.ac.uk, gye4@sheffield.ac.uk
@@ 2.3 Modality Bias
Prior work can be grouped into two directions. Diagnostic works primarily focus on revealing bias. For example, gradient-based analyses Kwon et al. (2025) and the attention sink studies Xiao et al. (2023) provide valuable insights into how linguistic shortcuts dominate inference but provide little in terms of mitigation. In contrast, mitigation methods typically rely on either extra branches Cadene et al. (2019); Clark et al. (2019), which introduce a language-only predictor to discount bias but double computation, or counterfactual data augmentations Chen et al. (2020), which enforce visual reliance through handcrafted or synthetic mismatched pairs. Although effective, both are costly or unscalable, limiting large-scale video understanding.
@@ 3.3 Language Advantage
Language can benefit TAL by providing high-level semantics difficult to infer from visual evidence alone. However, because TAL is fundamentally a vision-centric task, language can also become a shortcut: the model may reduce the training objective by exploiting frequent textual patterns rather than grounding its prediction in the video content. As a result, the multimodal loss alone is not a reliable signal of whether language is truly helpful.
To make this distinction explicit, we compare the task loss of a vision-only branch, denoted as $\mathcal{L}_{v}$, with that of the corresponding vision-language branch, denoted as $\mathcal{L}_{vl}$. Here, $\mathcal{L}_{v}$ reflects how well the task can be solved from visual evidence alone, while $\mathcal{L}_{vl}$ measures the loss after language is incorporated. Based on this comparison, we define Language Advantage (LA) as $A_{lang}=\mathcal{L}_{v}-\mathcal{L}_{vl}$.
This quantity measures the utility of language relative to vision. When language provides complementary semantics and improves prediction beyond the visual baseline, we have $\mathcal{L}_{vl}<\mathcal{L}_{v}$ and thus $A_{lang}>0$. When language introduces misleading priors or overrides visual evidence, $\mathcal{L}_{vl}$ can even exceed $\mathcal{L}_{v}$, yielding a small or negative $A_{lang}$ and indicating harmful modality bias.
This definition also explains why directly optimizing the multimodal branch can be problematic. In conventional VLMs, the model is trained only through $\mathcal{L}_{vl}$, so any shortcut-induced loss reduction would create a positive feedback loop that further increases reliance on language. In contrast, LA uses the vision-only branch as a stable reference, allowing us to determine whether language is providing genuine complementary information or merely acting as a shortcut.
Alternating-Epoch Estimation. A naive implementation would require an additional vision-only branch to compute $\mathcal{L}_{v}$ at every iteration. To avoid this overhead, we adopt a lightweight alternating-epoch estimation inspired by reinforcement learning Schulman et al. (2015). Training alternates between vision-only and vision-language epochs. Vision-only epochs compute class-wise average losses $\bar{\mathcal{L}}_{v}^{(c)}$, which then serve as a stable proxy for $\mathcal{L}_{v}$ during vision-language epochs. The language advantage for the $l^{\text{th}}$ frame is thus estimated as
@@ 6   Backpropagate vision-only task loss: $\mathcal{L}_{v}$;
Algorithm 1 summarizes our alternating-epoch estimation of LA. During each vision-only epoch, we compute the task loss using only visual features and accumulate the class-wise average loss $\bar{\mathcal{L}}_{v}^{(c)}$. In the subsequent vision-language epoch, this quantity serves as a lightweight proxy for the vision-only loss $\mathcal{L}_{v}$. In this way, LA is anchored to a vision-only baseline without requiring an additional vision branch at every iteration.
This estimation is computationally efficient for three reasons. First, the two training phases share the same backbone and differ only in whether the language model is activated, avoiding architectural duplication. Second, vision-only epochs skip the language model entirely and are therefore cheaper than standard multimodal updates, while still providing a principled reference for calibrating modality contributions. Third, because LA is derived directly from task losses rather than dataset-level bias statistics, it is naturally model- and dataset-agnostic, making the same estimation strategy applicable across different TAL benchmarks and VLM backbones.
@@ 3.4 Debiasing Feature Aggregation
To dynamically regulate the contribution of language, we derive a reliability weight $\lambda_{lang}$ from the predicted LA $\hat{A}_{lang}$, using it as a utility signal to determine whether language provides information beyond vision and should therefore participate in vision-grounded localization. To avoid assigning positive weights to negative advantages, we clip it and scale it into $[0,1)$ through sigmoid mapping:
@@
Rather than relying solely on language features for downstream tasks, we preserve the visual feature $F_{vis}$ and treat language features as a bounded residual refinement. Inspired by residual learning He et al. (2016), this design keeps vision dominant and allows language to complement it only when beneficial, thereby mitigating modality bias. As illustrated in Fig. 3(b), the final representations for classification and localization are obtained via element-wise addition:
@@ 4.2 Comparison with SoTA Methods
To examine robustness to language model capacity, we evaluate two InternVL2.5 Chen et al. (2024b) variants that use Qwen2.5-1.5B and 3B as their language backbones. InternVL3 is chosen for its strong alignment with visual downstream tasks. ActionVLM-S-3B achieves a 0.6% mAP improvement over the 1.5B counterpart on THUMOS14, suggesting that broader knowledge and richer pretraining better support TAL.
@@ 4.3 Analysis on Modality Bias
Measuring Modality Bias. To quantify modality bias at the performance level, we adopt Language Advantage on Performance (LAP), inspired by the diagnostic protocol of Cadene et al. (2019). Unlike the loss-based LA used during training, LAP evaluates how much prediction quality deteriorates when language priors conflict with the visual content. Specifically, we replace the aligned action descriptions with mismatched ones and measure the relative drop in average mAP. To create conflicts, we preferentially swap descriptions between semantically similar actions. When no clear counterpart exists, we randomly shuffle the description-action pairs. A larger LAP indicates that the model is more easily misled by text and therefore more biased toward linguistic cues, whereas a smaller LAP indicates stronger reliance on visual evidence and better robustness to misleading language.
We further examine two extreme hallucination patterns caused by modality bias. 1) Fixed outputs refer to nearly identical temporal segments repeatedly predicted within a sliding video window regardless of the input content, which we detect by checking whether segment intervals remain unchanged across consecutive steps. 2) Infinite listing refers to the model repeatedly emitting previously predicted segments without producing an End-of-Sequence token, which we detect by matching tail segments against earlier predictions. As shown in Table 2, QA-based VLMs exhibit both high LAP and non-trivial hallucination rates, whereas our FA formulation reduces LAP to 2.6 and eliminates both failure modes, indicating substantially stronger visual grounding.
@@ 4.4 Analysis on Language Advantage
To analyze how Language Advantage (LA) affects performance, we group action categories in THUMOS14 into easy, medium, and hard subsets based on the performance of the vision-only baseline (detailed in Appx. A.9). The hard subset includes Billiard and FrisbeeCatch. While multiple factors may contribute to their lower performance, both actions exhibit notable visual ambiguity: Billiard exhibits subtle and sparse motion cues, while FrisbeeCatch is easily confused with throwing actions and differs only in temporal pose order (Fig. 4).
@@ 5 Conclusion
In this study, we present ActionVLM, a vision-centric VLM for Temporal Action Localization (TAL) that explicitly addresses modality bias. Rather than allowing language to dominate prediction through shortcut correlations, our framework preserves visual grounding and invokes language only when it provides measurable benefit, through language advantage estimation and debiased residual aggregation. In this way, ActionVLM enables language to refine rather than override visual grounding. Extensive experiments show that this design not only alleviates modality bias and improves robustness under visual ambiguity, but also yields consistent gains across TAL benchmarks with negligible overhead. More broadly, our results suggest a practical multimodal alternative to purely visual scaling, pointing toward robust video understanding by better calibrating language priors rather than relying solely on ever larger visual data and models.
Limitations
@@
This work builds upon publicly available video datasets and does not involve private or personally identifiable data. However, vision-language models may inherit biases from pre-training corpora and textual supervision, which can propagate into temporal action localization. If misused in surveillance or monitoring settings, action localization systems could enable harmful profiling or intrusive behavioral analysis. In addition, biased descriptions or uneven data coverage may affect some activities, populations, or environments disproportionately when such systems are deployed in the real world. To mitigate these risks, we recommend dataset auditing, bias evaluation across activity and environment groups, and uncertainty reporting before any deployment. Our method is intended for academic research on robust and fair video understanding rather than high-stakes decision-making.
@@ A.2 Analysis on Efficiency
Table 6 compares ActionVLM with both conventional VLMs and strong vision-only baselines in terms of computational efficiency and accuracy on THUMOS14. We report average training and inference latency per video clip. The window size denotes the number of frames processed in a single temporal chunk, which affects both efficiency and temporal context coverage. We set a smaller window size for conventional VLMs to fit GPU memory, while ActionVLM and AdaTAD use longer windows to capture richer temporal dynamics following the original setting Liu et al. (2024).
@@ A.3 Additional Experiments on EPIC-Kitchens
To validate the generalizability of our model on diverse datasets, we further conduct evaluation on EPIC-Kitchens-100 Damen et al. (2020). EPIC-Kitchens-100 is a large-scale egocentric benchmark comprising 700 videos (about 100 hours) recorded in 45 kitchens, with fine-grained verb and noun annotations. Its domain specificity and egocentric action patterns differ substantially from the visual and linguistic distributions commonly seen in large-scale VLM pretraining corpora, making high absolute mAP particularly challenging Liu et al. (2024).
We use GPT-4o-mini to generate short descriptions for each class used in our EPIC-Kitchens-100 evaluation setup. This results in 1,283 description tokens for the 78 verb classes and 2,277 tokens for the 211 noun classes. Importantly, our runtime is dominated by video encoding rather than language modeling, so the text token budget is not the computational bottleneck.
@@ A.6 Histogram of Linguistically Preferred Boundaries
A.7 Model Sensitivity to Language Granularity
Action descriptions improve TAL by providing auxiliary semantic priors, but they also introduce an additional text-token budget. We further ask a practical question: how much description detail is actually needed? In particular, we study whether the full sentence descriptions used in ActionVLM can be replaced by cheaper alternatives without sacrificing accuracy.
We compare two budget-saving variants against the full setting. The first is keywords-only, where each class is represented by only three words summarizing the start, middle, and end motions. The second is merged descriptions, where semantically similar actions (e.g., LongJump&HighJump) share one coherent sentence. Table 10 reports the results.
Model #Token mAP
w/o description 0 70.7
keywords-only 219 71.0
merged descriptions 659 71.9
ActionVLM-S-3B (full) 742 72.0
Table 10: Sensitivity to language granularity on THUMOS14.
While keywords-only descriptions reduce the text budget substantially, they provide only marginal gains over removing descriptions, whereas merged descriptions nearly match the full model. This indicates that preserving coherent action semantics is more important than retaining every class-specific phrase. Since our efficiency analysis in Table 6 shows that our runtime is dominated by video encoding rather than language modeling, we use the full descriptions by default to achieve the best performance.
A.8 Analysis on Performance Gain
@@
A.9 Action Categories by Visual Ambiguity